OpenStudy (anonymous):

) If I place 3 moles of N2 and 4 moles of O2 in a 35 L container at a temperature of 250 C, what will the pressure of the resulting mixture of gases be?

OpenStudy (accessdenied):

The first key detail is that we have a mixture of gases. Dalton's Law applies here. P_total = the sum of the Partial Pressures of each gas as it would be alone. That leaves us to determine these two partial pressures: (a) 3 moles N2 in a 35 L container at 250 C (b) 4 moles O2 in a 35 L container at 250 C That is clear?

OpenStudy (anonymous):

yes

OpenStudy (accessdenied):

So we will need to then apply the ideal gas law equation because we have a volume, an amount (moles), and a temperature. We only need pressure. PV = n R T We would find P = (n R T) / V in both cases and add the two results. Conveniently, you are given the moles right off the bat so we don't need to worry about the compounds involved (we could use it to determine the moles from a mass otherwise).

OpenStudy (anonymous):

How would we do that

OpenStudy (accessdenied):

Convert the temperature to Kelvins first: T = 250 + 273 = 523 K In case (a), we have V = 35 L. T = 523 K. n = 3 moles N2. There are several R values and we have not specified which pressure unit to use. I usually remember kiloPascals (R = 8.31 kPa L/(mol K)) . If you feel more comfortable using another unit, it is an easy adjustment. Anyways we just plug in the values and solve for P: P (35 L) = (3 mol N2 ) (8.314 (kPa L)/(mol K) (523 K) P = (3)(8.314)(523) / (35) kPa and the same is done for 4 mol O2.

OpenStudy (anonymous):

so would it be P= (4) (8.314) (523)/ thats all i got

OpenStudy (accessdenied):

for 4 mol, that is close. Just missing the division by 35 L. (a) P = 3(8.314)(523)/ 35 kPa (b) P = 4(8.314)(523) /35 kPa So in total, you add these two values to get the total pressure. (in kPa)

OpenStudy (anonymous):

would it be about 869.7kpa

OpenStudy (accessdenied):

Looks good to me. :)

OpenStudy (anonymous):

thank you

OpenStudy (accessdenied):

you're welcome!